OpenStudy (anonymous):

if x=-1 is a root of x^3+2x^2-x-20 use synthetic division to factor the polynomial and it all real solution of the equation i think the answer is (x+2)(x+1)^2 =-2-1

OpenStudy (anonymous):

well did you do synthetic division?

OpenStudy (rational):

is that the complete problem ?

OpenStudy (rational):

i dont quite get what do u mean by x=-1 is a root x^3+2x^2-x-20 well x = -1 is NOT a root for that polynomial

OpenStudy (anonymous):

yes i did the synthetic division

OpenStudy (anonymous):

and what did you get?

OpenStudy (anonymous):

yes that the whole problem i dont get it either my teacher wrote it on the board and told us to do it

OpenStudy (rational):

another bad problem wid poor phrasing, you cant do it :/

OpenStudy (rational):

here are the roots of given polynomial : http://www.wolframalpha.com/input/?i=roots+of+x%5E3%2B2x%5E2-x-20

OpenStudy (rational):

-1 is not one of them

OpenStudy (rational):

so the input to problem itself is messed up

OpenStudy (anonymous):

alright i just wont do it then and hope the teacher will explain it . thank you for the help

OpenStudy (rational):

wait a sec, i think i found the mistake :)

OpenStudy (rational):

we can solve it if the polynomial is \( x^3+2x^2-x-2\) instead of \( x^3+2x^2-x-2\color{red}{0}\)

OpenStudy (rational):

then it factors to : \(x^2 + 2x^2 -x -2 = (x+1)(x-1)(x+2)\) http://www.wolframalpha.com/input/?i=factor+x%5E3%2B2x%5E2-x-2

OpenStudy (rational):

just check once if u have copied the polynomial correctly